OpenStudy (anonymous):

The formula A = 118e0.024t models the population of a particular city, in thousands, t years after 1998. When will the population of the city reach 140 thousand? Show your work.

OpenStudy (anonymous):

my answer 2005 what my teacher said Substitute the values given to find the final answer. t is years after 1998 so once you have t add this to the year 1998.

OpenStudy (anonymous):

thats what i thought

OpenStudy (phi):

\[ 114 e^{0.024t} = 140 \\ e^{0.024t} = \frac{140}{114} \\ 0.024 t = \ln\left(\frac{140}{114} \right) \\ t= \frac{\ln\left(\frac{140}{114} \right) }{0.024} \]

OpenStudy (madhu.mukherjee.946):

8.56

OpenStudy (phi):

t= 8.56 years add this to 1998

OpenStudy (phi):

you get 2006.56 which is the middle of 2006

OpenStudy (madhu.mukherjee.946):

yeah